consider a constant payment “interest only” mortgage loan. the loan balance is L with the annual interest rate of r for t years. calculate the monthly payment
Answer #1

Monthly Payment= (L * (r/12*100) * ( 1+ r/12*100)^12t / ((1+r/12*100)^12t -1)

Lr/1200 * ((1200+r)/1200)^12t / {[((1200+r)/1200)^12t] - 1}

If ((1200+r)/1200)^12t = M

MP = Lr/1200 * M/(M-1)